OpenStudy (anonymous):

If a car is speeding down a road at 40 miles/hour (mph ), how long is the stopping distance \texttipD 40 D 4 0 compared to the stopping distance \texttipD 25 D 2 5 if the driver were going at the posted speed limit of 25 mph ?

OpenStudy (mathmale):

Not your fault, Motlox, but your posting is quite difficult to decipher. You may have to use the Draw feature (below) to present that numerical info.

OpenStudy (ddcamp):

Both cars would have the same de-acceleration (of gμ, where μ is the coefficient of friction). We have the kinematic equation: \[v_f^2 = v_i^2+2ad\] The final velocity will be 0, so we can set the equations for each scenario equal to each other: \[40^2+2aD_{40} = 25^2 + 2aD_{25} \\ Solve\ for\ \frac{ D_{40} }{ D_{25} }.\]